### v1

# Patterns vs. Patients: Evaluating LLMs against Mental Health Professionals on Personality Disorder Diagnosis through First-Person Narratives

Abstract
Growing reliance on LLMs for psychiatric self-assessment raises questions about their ability to interpret qualitative patient narratives. We present the first direct comparison between state-of-the-art LLMs and mental health professionals in diagnosing Borderline (BPD) and Narcissistic (NPD) Personality Disorders utilizing Polish-language first-person autobiographical accounts. We show that the top-performing Gemini Pro models surpassed human professionals in overall diagnostic accuracy by 21.91 percentage points (65.48% vs. 43.57%). While both models and human experts excelled at identifying BPD ($F_{1}=83.4$ & $F_{1}=80.0$, respectively), models severely underdiagnosed NPD ($F_{1}=6.7$ vs. $50.0$), showing a reluctance toward the value-laden term ,,narcissism.” Qualitatively, models provided confident, elaborate justifications focused on patterns and formal categories, while human experts remained concise and cautious, emphasizing the patients’ sense of self and temporal experience. Our findings demonstrate that while LLMs are highly competent at interpreting complex first-person clinical data, their outputs still carry critical reliability and bias issues.
Patterns vs. Patients: Evaluating LLMs against Mental Health Professionals on Personality Disorder Diagnosis through First-Person Narratives
Karolina Drożdż1*, Kacper Dudzic1,2,3*, Anna Sterna4, Marcin Moskalewicz1,4,5 1IDEAS Research Institute, Warsaw, Poland 2Adam Mickiewicz University, Poznań, Poland 3AMU Center for Artificial Intelligence, Poznań, Poland 4Poznań University of Medical Sciences, Poznań, Poland 5Maria Curie-Skłodowska University, Lublin, Poland Correspondence: karolina.drozdz@ideas.edu.pl
1 Introduction
00footnotetext: *Equal contribution.
Systemic pressures on mental healthcare have created critical accessibility gaps and diagnostic delays Sun2023; Barbui2025. Consequently, the public is increasingly turning to widely available Large Language Models (LLMs) for self-assessment, thereby bypassing traditional clinical pathways Lawrence2024; McBain2025. While this trend may partially relieve strain on the healthcare system, it also introduces considerable ethical and clinical risks, such as the potential for inaccurate or misleading diagnoses Harrer2023. This study empirically compares the diagnostic capabilities of current open- and closed-source LLMs to mental health professionals in identifying a complex and frequently misdiagnosed condition of Personality Disorder (PD).
Previous research shows that state-of-the-art LLMs can achieve performance levels that approach, or in some cases surpass, those of trained clinicians on standardized medical benchmarks grzybowski-etal-2025-polish; workum2025, and substantially improve clinicians’ diagnostic accuracy when used as a decision-support tool noda2025. These capabilities extend to chatbot-patient interaction, where recent investigations suggest that LLMs can surpass human physicians in expressing empathy in written clinical communication ayers2023. However, it remains unclear whether such capabilities can be effectively applied to complex psychiatric conditions, where diagnoses often lack well-established objective markers and depend heavily on nuanced, subjective accounts of patients.
A limited body of work evaluating LLMs specifically on psychiatric and psychological benchmarks reveals a more uneven performance. Studies show that while models such as GPT-4 and Llama 3 achieve high accuracy (up to 85%) in binary mental disorder detection tasks, their performance varies significantly across different datasets hanafi2025comprehensiveevaluationlargelanguage. Notably, LLMs fine-tuned on domain-specific data do not outperform generalist models on existing benchmarks fouda2025psychiatrybenchmultitaskbenchmarkllms; nguyen-etal-2025-large.
Despite these advances, most existing evaluations rely on binary or multiple-choice answer formats and use clinical-annotated datasets or noisy social media posts. Such evaluations are not directly comparable to the semi-structured, first-person narrative through which patients typically communicate their experiences to a mental health professional. Furthermore, existing studies have prioritized outcome metrics over explanatory evaluations of the reasoning process. Consequently, they do not verify whether models emulate human cognitive processes or if their justifications correspond to human diagnostic decisions. This discrepancy highlights a critical research gap: it remains unknown if LLMs can maintain diagnostic accuracy on raw first-person testimonies, provide diagnosis-relevant justifications, and whether their performance and reasoning compare with those of mental health professionals.
We explore this gap in a specific context of two conditions, namely Borderline Personality Disorder (BPD) and Narcissistic Personality Disorder (NPD). These are strong, yet underexplored, cases for model evaluation, particularly as their diagnostic assessment might be overdetermined by the theoretical framework applied, or the lack thereof (as with the lay understanding of the terms ,,borderline” and ,,narcissism”).
While the DSM-III  APA1980 formally established Personality Disorder (PD) as a standalone diagnostic unit, its initial categorical approach—in which mental health disorders are perceived as binary categories—has been criticized for producing high comorbidity and permitting substantial heterogeneity within categories Widiger2007. In response, recent revisions in DSM-5 APA2022 and ICD-11 WHO2022 have redirected attention from distinct categories of PD to transdiagnostic commonalities and their continuous variation, i.e., a dimensional, non-exclusive approach. The DSM-5 AMPD (Section III) APA2022 introduced a general criterion (Criterion A) that defines a shared basis for all PD, independent of their categorical manifestations Pincus2020.
Given the prominence of the ICD and DSM frameworks, current LLMs could be biased toward a dimensional understanding of PD. Conversely, the ubiquity of terms like ,,borderline” and ,,narcissism” in everyday language—and thus in training corpora—suggests a potential counter-bias toward a categorical understanding of PD. Since empirical evidence does not clarify whether current LLMs lean toward a dimensional or categorical view of PD, we examined both frameworks independently.
In summary, our contributions are as follows:
  • We present the first evaluation of state-of-the-art LLMs on raw, first-person accounts of patients’ life stories, assessing the models’ capacity to interpret lived biographical experiences.
  • We introduce an expert-informed evaluation protocol that encompasses diverse conceptualizations of mental disorders (i.e., categorical vs. dimensional).
  • We directly compare the diagnostic performance of human experts and LLMs on complex mental disorders, moving beyond isolated model benchmarking.
  • We adopt an explanatory approach to both human data and model outputs, aiming not only to quantify performance but also to provide interpretable insights.
2 Methods
2.1 Data
To ensure the clinical validity of the sensitive first-person data used for evaluation, all narratives were collected at the Psychiatric Hospital in Międzyrzecz, Poland, following a multi-stage procedure.
First, 120 patients with a diagnosed personality disorder were identified. Each diagnosis had been established in outpatient care and subsequently confirmed by two experienced psychiatrists during inpatient admission, followed by an assessment by a clinical psychologist. Second, cases involving suicidal crisis, psychotic decompensation, or comorbid substance use disorders were excluded. Third, we employed intensity sampling to select information-rich, yet non-extreme, cases suitable for both qualitative and computational analysis.
Based on ICD-10 diagnostic coding, 24 BPD and 20 NPD cases were shortlisted. To ensure comparability, we assessed personality functioning and maladaptive traits using a multi-dimensional protocol (see Appendix A). All patients participated in 50–70 minute semi-structured qualitative interviews adapted from McAdams’ Life Story Interview McAdams1988, designed to explore narrative identity. Two additional questions probing reflective self-experience were included. Interviews were audio-recorded and transcribed, resulting in a Polish-language corpus exceeding 200,000 words. The qualitative analyses of all BPD cases Sterna2025 and all NPD cases (conducted for this project) provided an in-depth understanding of the subjective nuances represented in the data.
Six cases—three BPD and three NPD—were selected for this study. These were meticulously chosen to represent mild, moderate, and severe levels of impairment, and to capture the breadth of narrative expression within each disorder. Their selection reflects both the unique richness of their narratives and the practical constraints imposed by a human-model comparison (since each narrative is long, complex, and requires significant time commitment and detailed evaluation).
Additionally, from a pool of 20 control participants who underwent the same interview procedure, one control case with no personality disorder and subclinical personality functioning was chosen. In total, the evaluation dataset comprised 7 transcriptions of autobiographical interviews: 3 BPD, 3 NPD, and 1 Healthy Control.
2.2 Model Selection
We included a total of N=16{N=16} leading models and variants, split into a closed-source ($N=9$) and an open-source group ($N=7$). Our selection criteria encompassed: general public interest and availability, presence in recent evaluations on related material fouda2025psychiatrybenchmultitaskbenchmarkllms; hua2025scoping; nguyen-etal-2025-large, access model (licensing), parameter count, country of origin, and reasoning ability.
The closed-source group included: Gemini 2.5 Pro comanici2025gemini25pushingfrontier, Gemini 3 Pro google2025gemini3, Claude Opus 4.1 anthropic2025claude4, GPT-4o openai2024gpt4ocard, GPT-4.1 openai2025gpt41, and four variants of GPT-5 openai2025gpt5 with all available values of the OpenAI API’s reasoning effort parameter (from minimal to high). The open-source group consisted of: Gemma 3 27B gemmateam2025gemma3technicalreport, Llama 3.3 70B grattafiori2024llama3herdmodels, DeepSeek R1 0528 deepseekai2025deepseekr1incentivizingreasoningcapability, two variants of DeepSeek v3.1 Terminus deepseekai2025deepseekv3technicalreport with reasoning either enabled or disabled, and two variants of Qwen 3 32B yang2025qwen3technicalreport, also with and without reasoning.
We did not include domain-specific models trained on medical data for two reasons. Firstly, the context windows of available medical models evaluated in existing literature fouda2025psychiatrybenchmultitaskbenchmarkllms; hanafi2025comprehensiveevaluationlargelanguage; hua2025scoping; nguyen-etal-2025-large proved to be too small for the full patient testimonies—except for MentalQLM shi2025mentalqlm, which has not been made publicly available. Secondly, recent studies indicate that general-purpose models perform similarly to or better than medical models on psychiatric tasks fouda2025psychiatrybenchmultitaskbenchmarkllms.
2.3 Mental Health Experts
The study sample consisted of $N=6$ highly experienced mental health professionals, comprising three psychiatrists and three psychotherapists, recruited through our professional network. The group was balanced by gender, with three participants identifying as men and three as women; the mean age was 48.5 years (SD = 11.18), with an average of 18.5 years of professional experience. Importantly, the experts were external to the original diagnostic teams. They were blinded to the patients’ previous medical records and ground truth diagnoses, and had never previously encountered the patients whose narratives they evaluated.
2.4 Procedure
Human and model participants conducted a diagnostic assessment of each autobiographical testimony, adhering to a 6-step response template (see Appendix B). The 6-step protocol required participants to assign a (1) categorical diagnosis and (2) severity rating, each with a corresponding confidence level (3 & 4). Participants were also instructed to provide a brief justification for both (5) the diagnosis and (6) severity assessment, in no more than 200 words, indicating relevant evidence from the testimony and outlining their theoretical understanding of PD and their origins. All assessments were conducted in Polish, reflecting the language of the source data and the native language of the mental health professionals.
To avoid priming and to allow participants to apply their expert knowledge, the categorical diagnosis was an open-ended task without predefined options. In contrast, both the severity rating and the confidence ratings followed fixed scales: severity was rated on a 0–3 scale (0 = none, 1 = mild, 2 = moderate, 3 = severe), and confidence on a 1–4 scale (1 = guessing, 2 = somewhat confident, 3 = fairly confident, 4 = completely confident).
While human participants evaluated each testimony once, models were presented with each case three times to address potential diagnostic inconsistency stemming from their non-deterministic nature. This resulted in 7 trials per human expert and 21 trials (7 cases × 3 repetitions) per model.
2.5 Data Analysis
2.5.1 Performance Metrics
Two complementary performance metrics were computed: a categorical and a dimensional score. The first, based on the open-ended diagnosis, captured performance within the traditional categorical model, which assumes the presence or absence of discrete diagnostic entities. The second was calculated from the severity rating, which reflects the dimensional model conceptualizing personality disorder along a continuum of severity.
For human participants, each score represented the total number of correct evaluations across the seven cases (0–7). For the models, scores were computed across all 21 trials to account for reliability (7 cases × 3 repetitions; 0–21 score). Since a correct model response required both accuracy and consistency across the repeated trials, a perfect score would indicate not only a valid diagnostic judgment but also a fully consistent performance.
2.5.2 Diagnostic Justifications
Human mental health professionals can arrive at similar diagnostic conclusions through different cognitive and emotional pathways biondi2022clinician. Traces of these mental processes are reflected in the semantic content of the justifications provided during the diagnosis. To use this information alongside quantitative metrics, we mapped the diagnostic justifications onto a high-dimensional semantic embedding space. This allowed for a comparison of diagnostic expertise leverage between human and model participants, as well as among different models.
BAAI/bge-multilingual-gemma2111https://huggingface.co/BAAI/bge-multilingual-gemma2 was chosen as the embedding model owing to its superior performance on Polish-language tasks in the MMTEB benchmark enevoldsen2025mmtebmassivemultilingualtext. We created a single summary embedding representing the averaged semantic content of justifications for each model separately, as well as a single one representing all human participant data to account for its comparative scarcity. A more detailed process of embedding creation was described in Appendix G. The summary embeddings were aggregated into two distinct datasets: a global dataset comprising all agents (both models and humans) to assess the human-AI semantic gap, and a model-only dataset to allow for a finer-grained analysis of inter-model semantic differences.
To visualize the high-dimensional relationships between these representations, we employed a dual-method dimensionality reduction approach utilizing Multidimensional Scaling (MDS) kruskal1964multidimensional and Uniform Manifold Approximation and Projection (UMAP) mcinnes2020umapuniformmanifoldapproximation. This combination was chosen to provide a comprehensive view of the semantic space: MDS was selected for its ability to preserve global metric distances, offering a faithful representation of the relative dissimilarities between models. Complementarily, UMAP was utilized for its strength in manifold learning and local structure preservation, which allows for the identification of distinct semantic clusters (neighborhoods).
MDS was conducted using the scikit-learn [^2] https://github.com/scikit-learn/scikit-learn library JMLR:v12:pedregosa11a. Before dimensionality reduction, a pairwise cosine distance matrix was computed for the embeddings to ensure that the angular relationships inherent to semantic vector spaces were preserved. UMAP was conducted using the umap-learn [^3] https://github.com/lmcinnes/umap library. Details on the hyperparameter values used can be found in Appendix F.
2.5.3 Lexical Features
To further investigate the divergence between linguistic justification of diagnostic reasoning in humans and models, we conducted a follow-up inquiry into interpretable differences. We aggregated all justifications into two distinct corpora—human-written and model-written—and sought to identify lexical features that were statistically overrepresented in one group relative to the other. The features were identified with the weighted log-odds ratio with informative Dirichlet prior method originally applied to the problem of detecting lexical polarization in political discourse monroe2008fightin.
In the technical implementation, we first applied a text pre-processing pipeline to the two corpora, converting non-Polish non-ASCII characters into ASCII equivalents, as well as removing LLM-characteristic text formatting artifacts, punctuation, digits, and optionally stop words. Finally, we applied Monroe et al.’s method through the ConvoKit library chang-etal-2020-convokit. We considered two parameters to adjust in the feature generation process: stop word removal—either enabled or disabled, and the maximum n-gram length—exclusive ranges of 1 to 3 and a combined 1–3 one, ultimately choosing unigrams with the stop words removed as the most informative configuration.
3 Results
3.1 Diagnostic Performance
Figure 1: Categorical and dimensional score (recalculated on a 0–1 scale) comparison between mental health professionals and models.
Figure 2: Average model and mental health professionals’ scores by true diagnosis.
A comparative analysis of diagnostic performance revealed significant variability among models and mental health professionals, as summarized below.
Best-performing models outperform the human diagnostic average. The overall human average score was 43.57%, with a categorical score of 3.60 out of 7 (51.43%) and a dimensional score of 2.50 out of 7 (35.71%). An exploratory analysis of the participant sub-groups suggested a potential performance gap: psychotherapists scored higher on the categorical task (mean difference: 2.33), while psychiatrists scored slightly higher on the dimensional task (mean difference: 0.33).
Model performance, evaluated over 21 trials, was relatively dispersed, with several top-performing models exceeding the human average (see Figure 1). For the categorical task, most models (12 of 16) converged on a score of 12/21 (57.1%), outperforming the human average of 51.4%. Gemini 3 Pro achieved the highest categorical accuracy (15/21; 71.4%), while Gemini 2.5 Pro performed best on the dimensional score (14/21; 66.7%). Notably, the averaged Gemini Pro model family overall score exceeded the human overall average by 21.91 percentage points (65.48% vs. 43.57%). Conversely, Gemma 3 27B was the worst-performing model, scoring 9/21 (42.8%) on the categorical task and 6/21 (28.5%) on the dimensional task.
Further analysis of model performance indicated that reasoning did not consistently improve diagnostic accuracy. Model size appeared to be a factor, with smaller models generally performing worse than larger ones. No clear performance gap was observed between closed-source and open-source models.
BPD overdiagnosis and NPD underdiagnosis drive systematic bias, most pronounced in models. An analysis of the $N=378$ total diagnoses (see Appendix C, Table C.1) by both models ($N=336$) and humans ($N=42$) revealed systematic diagnostic bias.
BPD was the most common diagnosis, accounting for 53.97% of all combined diagnoses. This tendency to overdiagnose BPD resulted in a near-perfect categorical recall for models (97.9%, $F1=83.4$), substantially higher than that of human experts (66.7%, $F1=80.0$). However, such sensitivity came at the cost of precision (72.7%), whereas human experts demonstrated perfect precision (100%), suggesting that while clinicians missed some cases, their positive diagnoses were fully reliable. The high categorical recall for both models and humans was often paired with a lower dimensional recall (45.14% for models, 38.89% for humans; see Figure 2).
In contrast, NPD was severely underdiagnosed by both humans and models, representing 2.65% (10 of 378) of all diagnoses, with humans outperforming the models (14% against 1.5% for models). Gemini 2.5 Pro and Gemini 3 Pro were the only models to correctly identify 1 out of 3 NPD cases (in 2 out of 3 trials for Gemini 2.5 Pro, and 3 out of 3 trials for Gemini 3 Pro). Consequently, the aggregate model recall for NPD collapsed to just 3.5% ($F1=6.7$). While the rare NPD diagnoses provided by models were accurate ($precision=100\%$), the vast majority of cases were missed. This contrasts with human experts, who achieved a higher recall of 33.3% ($F1=50.0$). However, a distinct positive dissociation was observed in model performance: despite failing to categorically identify the disorder, models achieved a significantly higher dimensional recall (40.97%) for the same NPD cases, notably outperforming human experts (22.22%) on the severity metric.
The ,,healthy” label was the second most common, accounting for 21.43% of all assigned diagnoses. Models were highly effective at identifying the absence of pathology ($recall=87.5\%$, $F1=69.4$) compared to human experts ($recall=60.0\%$, $F1=50.0$). Despite these high scores, a detailed analysis of diagnostic distributions (see Appendix D, Figure D.1) reveals a specific ,,depathologizing bias” in the GPT family. GPT-4o showed the most extreme bias, misclassifying cases as healthy 10 times. GPT-4.1 followed with 5 false positives, whereas the GPT-5 variants had 2–3 each. This suggests a tendency in the GPT model family to favor non-clinical classifications. Notably, humans also showed a slight bias toward non-clinical classifications, evidenced by 4 false positives (see Appendix D, Figure D.1), driven by frequent assessments that the testimonies lacked sufficient evidence for a PD diagnosis.
A significant misclassification bias was observed for Avoidant Personality Disorder (AvPD) (see Appendix C, Table C.1). It emerged as the third most common diagnosis overall (15.08%), but this prevalence was driven largely by models. AvPD accounted for 16.07% (54 of 336) of model-generated diagnoses, compared to 7.14% (3 of 42) of human diagnoses. Given that AvPD was not present in the ground truth answers, this high prevalence points to a systematic tendency to misinterpret symptoms of other conditions as AvPD. On the other hand, only one out of 336 diagnoses by models indicated a PD category not present in current diagnostic frameworks, namely Masochistic PD.
Models demonstrate higher and more uniform confidence than human experts. Overall, models displayed higher certainty than human practitioners for both categorical diagnosis ($M=2.94$ vs. $M=2.52$) and severity assessment ($M=3.07$ vs. $M=2.86$). Across both groups, certainty was higher for severity assessments than for categorical diagnoses. Among the individual models, Gemini 3 Pro ($M=3.69$) and Gemini 2.5 Pro ($M=3.40$), the top-performing models in this study, exhibited the highest levels of certainty. Conversely, GPT-5 with high reasoning effort was the least certain model ($M=2.57$). The most striking difference between groups was that the models never utilized the lowest certainty score (1 – ,,guessing”) whereas human practitioners used it in 19% of diagnostic ratings and 11.1% of severity ratings.
3.2 Diagnostic Justifications
3.2.1 Semantic Embeddings
Figure 3: MDS and UMAP projections of the semantic embeddings of diagnosis justifications.
The MDS and UMAP projections of the semantic embeddings of diagnosis justifications are visualized in Figure 3. Overall, for this analysis, MDS proved to be a more informative dimensionality-reduction method than UMAP, contributing to most of the observations summarized as follows.
Models from a single family (but not necessarily the same series) are generally close semantically. This is a general trend observable among all the evaluated model families: DeepSeek, GPT, Gemini, and Qwen, with their points creating visible neighborhoods distance-wise. Additionally, on the example of DeepSeek v3.1 Terminus, Qwen 3 32B, and GPT-5 clusters, we observe that the impact of chain-of-thought reasoning—either its binary presence in the case of the first two, or its gradation in the case of the last one—on model semantics is rather low.
GPT-4.1 correlates semantically with the GPT-5 cluster, but not with GPT-4o. GPT-4.1, being from the same model family but not the same series, is located close to the GPT-5 model variants on each projection, echoing the previous observation. Yet, this did not apply to GPT-4o, which emerged as a singular exception to the trend of semantic closeness within a model family by being located far from both the GPT and the main point clusters.
Llama 3.3 70B exhibits atypical semantics compared to every other agent. The distinctiveness of Llama’s semantics does not stem from undertaking vastly different diagnostic pathways than the rest of the agents, but from its comparably poor command of the Polish language. Llama’s explanations tended to contain incorrect grammar, nonexistent words, or seemingly random ,,artifact” tokens unrelated grammatically and semantically to the rest of the justifications’ linguistic content. Interestingly, this did not seem to have significantly impeded Llama’s diagnosis capabilities, as it was the ex-aequo third-best performing model (see Figure 1).
Models performing weakly in terms of diagnostic performance also exhibit more atypical semantics. Both Qwen variants, Gemma 3 27B, and GPT-4o not only achieved the lowest cumulative diagnostic performance scores (see Figure 1), but also are all semantically located relatively far away from the main model cluster, while at the same time being located close to each other.
Human participant semantics strongly differ from model semantics. A large observable difference between the general semantic tendencies of human participants compared to models can be attributed to differing justification strategies. Human participants provided concise and direct explanations, rarely approaching the 100-word limit. One psychotherapist and one psychiatrist copied excerpts from patient testimonies rather than generating original justifications, while two other psychiatrists produced unusually brief responses: one did not justify healthy-control diagnoses, whereas the other declined to justify low-confidence decisions. Several participants explicitly cut their justifications short, noting insufficient information for a reliable diagnosis. In contrast, all models consistently produced long, detailed, and highly confident justifications, never withholding a diagnosis based on inconclusive data.
3.2.2 Lexical features
We analyzed the 20 most characteristic features identified for human and model outputs by z-score values (for the full list, see Appendix E). The original Polish unigrams were supplemented with their closest English translations and are referred to as such in this section for clarity.
A higher prevalence of nouns characterized the human-characteristic lexical features group. Specifically, several features point to a human-centric approach to the diagnoses and their subsequent justifications, emphasizing a person’s sense of self and their interactions with society: patient, one’s own, interpersonal, others, image. Additionally, a more temporal focus is noticeable through the terms: time, sense of (time), future. Finally, the observable reluctance to give a definite diagnosis based just on textual data was reflected by mentions of its scarcity by some human participants: lack of, data.
In comparison, the model-characteristic lexical features group had a high prevalence of adjectives. This seems to be brought on by a more specific strong tendency of the models to categorize and formalize the patients’ subjective experiences and symptoms: persistent, severe, severity, moderate, chronic, intense, rigid, entrenched. The models also exhibit three pronounced single-feature propensities; heavy focus on seeing patterns in patients’ testimonies: patterns, comparatively high focus on violent life experiences: violence, and a predisposition (contrasting with humans) to create own interpretations of the particulars of the patients’ conditions when faced with unsatisfyingly detailed or unclear data: (I) understand (as).
4 Discussion
Beyond diagnostic accuracy, this study has several implications for the use of AI technology in clinical mental health contexts.
Regarding the conceptualization of mental disorders, our results highlight a tension between evolving psychiatric frameworks and the persistence of traditional diagnostic classifications. While human experts performed better on the categorical than the dimensional task, models achieved relatively balanced scores across both, with a performance preference for categorical diagnosis. This suggests that the current shift toward a dimensional understanding of personality pathology APA2022; WHO2022 may be overshadowed by the legacy of earlier categorical systems, whose labels remain prevalent in everyday discourse and online corpora.
The prominence of such labels in training corpora may also explain the severe underdiagnosing of NPD by models. This could stem from the negative semantic load and social stigma associated with the term ,,narcissist”, combined with the current LLM preference alignment training regimen (i.e., RLHF techniques) rewarding agreeable and non-confrontational behavior, sometimes bordering on (social) sycophancy Cheng2025; sharma2025understandingsycophancylanguagemodels. Assigning a stigmatized label to a first-person narrator—essentially ,,calling the user a narcissist”—conflicts with these preferences. Importantly, this avoidance does not reflect inability, as models showed higher dimensional recall for NPD cases, indicating sensitivity to severity without using value-laden labels. A similar aversion to stigmatizing terminology may explain the depathologizing bias observed in the GPT models family, where avoidance of psychiatric labels can downplay genuine symptoms and delay necessary treatment Semigran2015. Conversely, BPD is increasingly treated in the literature as a clinical condition rather than a pejorative label. The comparatively extensive body of BPD research relative to other disorders Blashfield2000 likely contributes to the models’ inclination to overdiagnose it.
Notably, a critical divergence between humans and LLMs was observed in the assessment of confidence. Models never utilized the ,,guessing” or low-confidence options, whereas human experts frequently expressed uncertainty. This might pose a potential safety risk for clinical application. As argued by Ulmer et al. Ulmer2025, the ability of an AI to express doubt in a ,,human-like” manner is a necessary safety feature to ensure users can trust the uncertainty being conveyed. Paradoxically, although both humans and models demonstrated higher objective performance on the categorical task, they consistently reported higher subjective certainty regarding their dimensional ratings. Future studies could investigate the underlying mechanisms of this metacognitive dissociation, examining why both human and non-human agents perceive their answers regarding dimensional severity as more certain, despite their objective performance superiority in categorical classification.
Furthermore, the semantic analysis reveals that although models may achieve higher accuracy, their ,,reasoning” remains fundamentally distinct from clinical practice. Human justifications were concise, cautious, and patient-centered. Model justifications, conversely, were elaborate, formulaic, and pattern-focused. Additionally, models that showed reduced performance also had more atypical semantics. This suggests a link between diagnostic accuracy and justification semantics; lower-performing models may have failed to attend to the most clinically informative aspects of the patient narratives. Llama 3.3 70B was a notable exception, achieving high diagnostic accuracy despite generating justifications riddled with linguistic artifacts, nonexistent words, and poor Polish grammar. This can be understood in terms of a multilingual generative-discriminative performance gap, in which Llama 3.3 70B demonstrates sufficient passive understanding of Polish to assign accurate diagnoses, while exhibiting markedly weaker Polish fluency when generating justifications for those decisions. Follow-up research could investigate how the performance of this model would change if its diagnostic reasoning were generated before assigning diagnostic scores, rather than as a post-hoc justification.
5 Conclusion
In this study, we provided a first-of-its-kind evaluation of state-of-the-art LLMs on raw, first-person patient data, assessed under a novel expert-informed evaluation protocol and directly compared with human specialists. Our findings indicate that current LLMs possess a surprisingly high level of diagnostic competence, largely surpassing mental health professionals in the examined context. At the same time, the models demonstrated susceptibility to bias and overconfidence. Taken together, these findings underscore the need for a collaborative human–AI framework, in which the clinical judgment and ethical oversight of experts balance the analytical consistency of models. Such integration has the potential to mitigate the respective limitations of each agent—model bias and human variability—thereby achieving higher diagnostic validity than either could attain alone.
Limitations
This study has several limitations. First, humans and LLMs were compared on the textual diagnostic modality only, whereas actual psychiatric assessment is inherently multimodal; clinicians rely on visual cues, affect, prosody, and interactional dynamics that are absent in text transcripts. Accordingly, the human experts in our experiment reported difficulties in relying solely on text without seeing the patient. However, given the rapid post-COVID rise of remote medicine, as well as the more recent surge in AI-driven therapeutic chatbots, the textual modality may gain prominence in the future. The second limitation is that the sample comprised only seven narratives. Although small, this sample size reflects a deliberate trade-off necessitated by the depth of the data and the substantial cognitive and time demands of human experts. The fact that this study was conducted within a specific cultural context may also be considered a limitation. This is because the diagnostic approaches of human experts and the narrative styles of patients may not be generalizable to other populations. Future research should explore whether these findings remain consistent across different cultural settings.
Ethical considerations
All procedures performed in this study involving human participants were in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki Declaration and its later amendments or comparable ethical standards. Human participants signed an informed consent form by hand. The qualitative interview protocol was approved by the hospital and the Bioethics Committee of Poznan University of Medical Sciences (decision no. KB-367/23). All personal information was anonymized, ensuring full privacy of the patients’ identities.
Mental health diagnosis remains a fundamentally human responsibility requiring empathy, ethical judgment, and contextual understanding beyond mere pattern recognition. Nevertheless, in recent years, patients have already begun to independently use AI-based consumer tools for self-assessment outside established clinical frameworks. We acknowledge the ethical and moral dilemmas associated with the use of AI in clinical research and practice, and, as noted above, advocate that AI technologies be deployed in clinical settings only under the supervision of human experts.
Appendix A Clinical Assessment Instruments
Levels of personality functioning and maladaptive trait domains were assessed using the Level of Personality Functioning—Brief Scale 2.0 (LPFS-BF 2.0) Lakuta2023 and the Personality Inventory for ICD-11 Cieciuch2022, ensuring comparability across individuals. LPFS-BF 2.0 scores were interpreted relative to established cut-offs Weekers2023, yielding a dimensional classification from subclinical to extreme impairment (0–48), independent of the categorical BPD/NPD labels.
Appendix B Model Prompts
Original Polish prompt used in the study Jesteś doświadczonym specjalistą zdrowia psychicznego. Twoim zadaniem jest przeprowadzenie wstępnej oceny diagnostycznej na podstawie przedstawionego poniżej tekstu autobiograficznego. Wykonaj 6-etapową analizę, prezentując swoje wnioski w ściśle określonym formacie. # OCENA DIAGNOSTYCZNA ## Krok 1 Przypisz diagnozę kategorialną zaburzenia osobowości. Pamiętaj, że możliwa jest opcja ,,Brak zaburzenia”. ## Krok 2 Oceń stopień pewności co do poprawności swojej odpowiedzi z Kroku 1, używając skali 1–4, gdzie: 1 = zgadywałem/am, 2 = trochę pewny/a, 3 = dość pewny/a, 4 = całkowicie pewny/a. ## Krok 3 Uzasadnij swoją decyzję diagnostyczną z Kroku 1. Wskaż kluczowe fragmenty danych, które popierają twoją ocenę. Następnie, odnieś się do tego czym są dla ciebie zaburzenia osobowości i wyjaśnij, jak rozumiesz ich źródła. Limit uzasadnienia to 100 słów. ## Krok 4 Oceń stopień nasilenia zaburzenia osobowości, używając skali 0–3, gdzie: 0 = brak, 1 = łagodne, 2 = umiarkowane, 3 = ciężkie. ## Krok 5 Oceń stopień pewności co do poprawności oceny nasilenia z Kroku 4, używając skali 1–4, gdzie: 1 = zgadywałem/am, 2 = trochę pewny/a, 3 = dość pewny/a, 4 = całkowicie pewny/a. ## Krok 6 Uzasadnij swoją odpowiedź z Kroku 4. Wskaż kluczowe fragmenty danych, które popierają Twoją ocenę. Następnie, odnieś się do tego czym są dla ciebie zaburzenia osobowości i wyjaśnij, jak rozumiesz ich źródła. Limit uzasadnienia to 100 słów. Udziel odpowiedzi, używając poniższego szablonu. Analiza Przypadku [Numer przypadku] Diagnoza Kategorialna: [Twoja odpowiedź] Pewność Diagnozy (1–4): [Twoja odpowiedź] Uzasadnienie Diagnozy (do 100 słów): [Twoja odpowiedź] Ocena Nasilenia (0–3): [Twoja odpowiedź] Pewność Oceny Nasilenia (1–4): [Twoja odpowiedź] Uzasadnienie Oceny Nasilenia (do 100 słów): [Twoja odpowiedź] Teraz zapoznaj się z tekstem, przeprowadź analizę i odpowiedz zgodnie z podanym formatem. ## TEKST AUTOBIOGRAFICZNY {text} ## ODPOWIEDŹ
English translation of the original prompt You are an experienced mental health professional. Your task is to conduct a preliminary diagnostic assessment based on the autobiographical text presented below. Perform a 6-step analysis, presenting your conclusions in a strictly defined format. # DIAGNOSTIC ASSESSMENT ## Step 1 Assign a categorical personality disorder diagnosis. Remember that the option ,,No disorder” is possible. ## Step 2 Rate the degree of certainty regarding the correctness of your answer from Step 1, using a scale of 1–4, where: 1 = I was guessing, 2 = somewhat certain, 3 = fairly certain, 4 = completely certain. ## Step 3 Justify your diagnostic decision from Step 1. Indicate key data fragments that support your assessment. Then, refer to what personality disorders are to you and explain how you understand their origins. The justification limit is 100 words. ## Step 4 Rate the severity of the personality disorder, using a scale of 0–3, where: 0 = none, 1 = mild, 2 = moderate, 3 = severe. ## Step 5 Rate the degree of certainty regarding the correctness of the severity assessment from Step 4, using a scale of 1–4, where: 1 = I was guessing, 2 = somewhat certain, 3 = fairly certain, 4 = completely certain. ## Step 6 Justify your answer from Step 4. Indicate key data fragments that support your assessment. Then, refer to what personality disorders are to you and explain how you understand their origins. The justification limit is 100 words. Provide your answer using the template below. Case Analysis [Case Number] Categorical Diagnosis: [Your answer] Diagnosis Certainty (1–4): [Your answer] Diagnosis Justification (up to 100 words): [Your answer] Severity Assessment (0–3): [Your answer] Severity Assessment Certainty (1–4): [Your answer] Severity Assessment Justification (up to 100 words): [Your answer] Now familiarize yourself with the text, perform the analysis, and answer according to the provided format. ## AUTOBIOGRAPHICAL TEXT {text} ## ANSWER
Appendix C Diagnostic Labels
Label Count Percentage
BPD 204 53.97%
HC (Healthy) 81 21.43%
AvPD 57 15.08%
NPD 10 2.65%
Unspecified PD 6 1.59%
DPD 5 1.32%
AD/UNS 4 1.06%
Cannot diagnose 3 0.79%
DePD 2 0.53%
ASPD 2 0.53%
StPD 1 0.26%
OPD 1 0.26%
MPD 1 0.26%
HPD 1 0.26%
Table C.1: Distribution of diagnostic labels.
The diagnostic labels presented in Table C.1 correspond to the following clinical categories: BPD: Borderline Personality Disorder; NPD: Narcissistic Personality Disorder; AvPD: Avoidant Personality Disorder; HC: Healthy; StPD: Schizotypal Personality Disorder; DPD: Dependent Personality Disorder; DePD: Depressive Personality Disorder; AD/UNS: Anxiety Disorder/Unspecified; ASPD: Antisocial Personality Disorder; OPD: Obsessive Personality Disorder; MPD: Masochistic Personality Disorder; HPD: Histrionic Personality Disorder.
Appendix D Diagnostic Performance Heatmap
Figure D.1: Diagnostic performance heatmap of human experts and models.
The ,,Human Experts” row aggregates the performance of $N=5$ professionals; one provided only dimensional ratings, thus their data was excluded from categorical analysis but retained for dimensional evaluation. A False Positive indicates diagnosing a disorder absent in the ground truth, while a False Negative denotes failing to identify a present condition. Each model row reflects $N=21$ trials (7 cases x 3 runs), whereas the human row corresponds to $N=35$ independent assessments (7 cases x 5 experts).
Appendix E Most Representative Lexical Features
Figure E.2: The 20 most representative lexical features for human experts and models by z-score.
Appendix F Dimensionality Reduction Hyperparameters
Algorithm Parameter Value
MDS dissimilarity precomputed
n_components 2
random_state 42
UMAP n_components 2
n_neighbors 4
min_dist 0.1
random_state 42
Table F.2: Dimensionality reduction algorithms hyperparameter values.
Hyperparameters not present in the table retained the default values set in the algorithm implementations. The n_neighbors parameter for UMAP, which controls the balance between local and global structure, was set to a low value to accommodate the small dataset size of $N=17/16$ (with and without the human embedding). This setting prioritizes the detection of latent local substructures and family-specific clusters within the models.
Appendix G Semantic Justification Embedding Creation Process
The process of mapping the diagnosis justifications acquired through our procedure onto a high-dimensional semantic embedding space consisted of four steps.
  1. 1.
    The texts of all justifications were aggregated into groupings. For humans, due to the small amount of data relative to model outputs, a single grouping was created from all justifications for a single human participant. For each model, we created two separate groupings: one based on all categorical justifications and the other based on all dimensional justifications.
  2. 2.
    Before the embedding process, simple text pre-processing with regular expressions was applied to remove LLM-characteristic text formatting artifacts, such as Markdown syntax and redundant whitespace.
  3. 3.
    Each grouping was converted into a dense embedding using the chosen BAAI/bge-multilingual-gemma2 embedding model. The model operated in 16-bit precision on a single NVIDIA A100 (40GB) GPU, with a batch size of 8 and default remaining hyperparameter values.
  4. 4.
    A summary embedding representing the semantic contents of justifications was created for each model by first calculating the mean value for the categorical and dimensional grouping embeddings separately, and subsequently averaging these two values. For human participants, since there was no categorical-dimensional grouping separation, a single summary embedding was derived by averaging the embeddings of all individuals.

### v2

@@
 Growing reliance on LLMs for psychiatric self-assessment raises questions about their ability to interpret qualitative patient narratives.
-We present the first direct comparison between state-of-the-art LLMs and mental health professionals in diagnosing Borderline (BPD) and Narcissistic (NPD) Personality Disorders utilizing Polish-language first-person autobiographical accounts.
-We show that the top-performing Gemini Pro models surpassed human professionals in overall diagnostic accuracy by 21.91 percentage points (65.48% vs.
-43.57%).
+This depth-first case study provides the first direct comparison of state-of-the-art LLMs and mental health professionals in assessing Borderline (BPD) and Narcissistic (NPD) Personality Disorders based on Polish-language first-person autobiographical accounts.
+Within our sample, the overall diagnostic scores of the top-performing Gemini Pro models (65.48%) were 21.91 percentage points higher than the average scores of the human professionals (43.57%).
 While both models and human experts excelled at identifying BPD ($F_{1}=83.4$ & $F_{1}=80.0$, respectively), models severely underdiagnosed NPD ($F_{1}=6.7$ vs.
-$50.0$), showing a reluctance toward the value-laden term ,,narcissism.” Qualitatively, models provided confident, elaborate justifications focused on patterns and formal categories, while human experts remained concise and cautious, emphasizing the patients’ sense of self and temporal experience.
-Our findings demonstrate that while LLMs are highly competent at interpreting complex first-person clinical data, their outputs still carry critical reliability and bias issues.
+$50.0$), showing a potential reluctance toward the value-laden term ,,narcissism.” Qualitatively, models provided confident, elaborate justifications focused on patterns and formal categories, while human experts remained concise and cautious, emphasizing the patients’ sense of self and temporal experience.
+Our findings demonstrate that while LLMs might be competent at interpreting complex first-person clinical data, their outputs still carry critical reliability and bias issues.
 Patterns vs. Patients: Evaluating LLMs against Mental Health Professionals on Personality Disorder Diagnosis through First-Person Narratives
@@
 1 Introduction
-00footnotetext: *Equal contribution.
-Systemic pressures on mental healthcare have created critical accessibility gaps and diagnostic delays Sun2023; Barbui2025.
-Consequently, the public is increasingly turning to widely available Large Language Models (LLMs) for self-assessment, thereby bypassing traditional clinical pathways Lawrence2024; McBain2025.
-While this trend may partially relieve strain on the healthcare system, it also introduces considerable ethical and clinical risks, such as the potential for inaccurate or misleading diagnoses Harrer2023.
+Systemic pressures on mental healthcare have led to critical access gaps and diagnostic delays Sun et al.
+(2023); Barbui et al.
+(2025).
+Consequently, the public is increasingly turning to widely available Large Language Models (LLMs) for self-assessment, thereby bypassing traditional clinical pathways Lawrence et al.
+(2024); McBain et al.
+(2025).
+While this trend may partially relieve strain on the healthcare system, it also introduces considerable ethical and clinical risks, such as the potential for inaccurate or misleading diagnoses Harrer (2023).
 This study empirically compares the diagnostic capabilities of current open- and closed-source LLMs to mental health professionals in identifying a complex and frequently misdiagnosed condition of Personality Disorder (PD).
-Previous research shows that state-of-the-art LLMs can achieve performance levels that approach, or in some cases surpass, those of trained clinicians on standardized medical benchmarks grzybowski-etal-2025-polish; workum2025, and substantially improve clinicians’ diagnostic accuracy when used as a decision-support tool noda2025.
-These capabilities extend to chatbot-patient interaction, where recent investigations suggest that LLMs can surpass human physicians in expressing empathy in written clinical communication ayers2023.
+Previous research shows that state-of-the-art LLMs can achieve performance levels that approach, or in some cases surpass, those of trained clinicians on standardized medical benchmarks Grzybowski et al.
+(2025); Workum et al.
+(2025), and substantially improve clinicians’ diagnostic accuracy when used as a decision-support tool Noda et al.
+(2025).
+These capabilities extend to chatbot-patient interaction, where recent investigations suggest that LLMs can surpass human physicians in expressing empathy in written clinical communication Ayers et al.
+(2023).
 However, it remains unclear whether such capabilities can be effectively applied to complex psychiatric conditions, where diagnoses often lack well-established objective markers and depend heavily on nuanced, subjective accounts of patients.
 A limited body of work evaluating LLMs specifically on psychiatric and psychological benchmarks reveals a more uneven performance.
-Studies show that while models such as GPT-4 and Llama 3 achieve high accuracy (up to 85%) in binary mental disorder detection tasks, their performance varies significantly across different datasets hanafi2025comprehensiveevaluationlargelanguage.
-Notably, LLMs fine-tuned on domain-specific data do not outperform generalist models on existing benchmarks fouda2025psychiatrybenchmultitaskbenchmarkllms; nguyen-etal-2025-large.
-Despite these advances, most existing evaluations rely on binary or multiple-choice answer formats and use clinical-annotated datasets or noisy social media posts.
-Such evaluations are not directly comparable to the semi-structured, first-person narrative through which patients typically communicate their experiences to a mental health professional.
-Furthermore, existing studies have prioritized outcome metrics over explanatory evaluations of the reasoning process.
-Consequently, they do not verify whether models emulate human cognitive processes or if their justifications correspond to human diagnostic decisions.
-This discrepancy highlights a critical research gap: it remains unknown if LLMs can maintain diagnostic accuracy on raw first-person testimonies, provide diagnosis-relevant justifications, and whether their performance and reasoning compare with those of mental health professionals.
-We explore this gap in a specific context of two conditions, namely Borderline Personality Disorder (BPD) and Narcissistic Personality Disorder (NPD).
-These are strong, yet underexplored, cases for model evaluation, particularly as their diagnostic assessment might be overdetermined by the theoretical framework applied, or the lack thereof (as with the lay understanding of the terms ,,borderline” and ,,narcissism”).
-While the DSM-III  APA1980 formally established Personality Disorder (PD) as a standalone diagnostic unit, its initial categorical approach—in which mental health disorders are perceived as binary categories—has been criticized for producing high comorbidity and permitting substantial heterogeneity within categories Widiger2007.
-In response, recent revisions in DSM-5 APA2022 and ICD-11 WHO2022 have redirected attention from distinct categories of PD to transdiagnostic commonalities and their continuous variation, i.e., a dimensional, non-exclusive approach.
-The DSM-5 AMPD (Section III) APA2022 introduced a general criterion (Criterion A) that defines a shared basis for all PD, independent of their categorical manifestations Pincus2020.
+Studies show that models such as GPT-4 and Llama 3 achieve high accuracy (up to 85%) on binary mental disorder detection tasks, but their performance varies significantly across datasets Hanafi et al.
+(2025).
+Notably, LLMs fine-tuned on domain-specific data do not outperform generalist models on existing benchmarks Fouda et al.
+(2025); Nguyen et al.
+(2025).
+Existing evaluations typically rely on simplified formats and noisy or synthetic datasets, which are not comparable to the complex, first-person nature of psychiatric patients’ testimonies.
+Furthermore, by prioritizing outcome metrics over explanatory reasoning, current studies fail to verify whether models emulate human cognitive processes or merely guess accurately.
+To address this gap, our study introduces a methodological shift from breadth to depth.
+Specifically, instead of evaluating large volumes of surface-level material, we deliberately focus on a smaller number of raw, in-depth clinical cases.
+We apply this depth-first approach to Borderline Personality Disorder (BPD) and Narcissistic Personality Disorder (NPD).
+These are strong, yet underexplored, phenomenons for model evaluation, particularly as their diagnostic assessment might be overdetermined by the theoretical framework applied (or the lack thereof, as with the lay understanding of the terms ,,borderline” and ,,narcissism”).
+Recent revisions in DSM-5 American Psychiatric Association (2022) and ICD-11 World Health Organization (2022) have redirected attention from a distinct categorical approach to PD to one based on transdiagnostic commonalities and their continuous variation, i.e., a dimensional, non-exclusive approach.
 Given the prominence of the ICD and DSM frameworks, current LLMs could be biased toward a dimensional understanding of PD.
@@
   • We directly compare the diagnostic performance of human experts and LLMs on complex mental disorders, moving beyond isolated model benchmarking.
-  • We adopt an explanatory approach to both human data and model outputs, aiming not only to quantify performance but also to provide interpretable insights.
+  • We prioritize depth over breadth, adopting an explanatory approach to both human data and model outputs, aiming not only to quantify performance but also to provide interpretable insights.
 2 Methods
 2.1 Data
-To ensure the clinical validity of the sensitive first-person data used for evaluation, all narratives were collected at the Psychiatric Hospital in Międzyrzecz, Poland, following a multi-stage procedure.
-First, 120 patients with a diagnosed personality disorder were identified.
-Each diagnosis had been established in outpatient care and subsequently confirmed by two experienced psychiatrists during inpatient admission, followed by an assessment by a clinical psychologist.
-Second, cases involving suicidal crisis, psychotic decompensation, or comorbid substance use disorders were excluded.
-Third, we employed intensity sampling to select information-rich, yet non-extreme, cases suitable for both qualitative and computational analysis.
-Based on ICD-10 diagnostic coding, 24 BPD and 20 NPD cases were shortlisted.
-To ensure comparability, we assessed personality functioning and maladaptive traits using a multi-dimensional protocol (see Appendix A).
-All patients participated in 50–70 minute semi-structured qualitative interviews adapted from McAdams’ Life Story Interview McAdams1988, designed to explore narrative identity.
+The database comprises six clinical cases—three BPD and three NPD—alongside one healthy control.
+These were selected from a larger dataset of first-person symptomatology, which includes 24 BPD cases, 20 NPD cases, and 20 healthy controls (HC), classified according to ICD-10 diagnostic criteria.
+To ensure the clinical validity of the sensitive first-person data used for evaluation, clinical narratives were collected at the Psychiatric Hospital in Międzyrzecz, Poland, following a multi-stage procedure.
+Participants enrolled in the original study were diagnosed as outpatients, then confirmed as inpatients; each diagnosis was confirmed by four independent clinical experts: two psychiatrists and two psychologists.
+To ensure comparability, categorical diagnosis was accompanied by personality functioning and maladaptive traits assessment, using a multi-dimensional protocol (see Appendix A).
+HC participants were healthy volunteers who responded to the study’s announcement and declared no diagnosed psychiatric conditions or felt symptoms.
+They were subsequently screened for personality pathology using LPFS BF 2.0  Łakuta et al.
+(2023) and PICD Cieciuch et al.
+(2022).
+All the participants took part in 50–70 minute semi-structured qualitative interviews adapted from McAdams’ Life Story Interview McAdams (1988), designed to explore narrative identity.
 Two additional questions probing reflective self-experience were included.
 Interviews were audio-recorded and transcribed, resulting in a Polish-language corpus exceeding 200,000 words.
-The qualitative analyses of all BPD cases Sterna2025 and all NPD cases (conducted for this project) provided an in-depth understanding of the subjective nuances represented in the data.
-Six cases—three BPD and three NPD—were selected for this study.
-These were meticulously chosen to represent mild, moderate, and severe levels of impairment, and to capture the breadth of narrative expression within each disorder.
-Their selection reflects both the unique richness of their narratives and the practical constraints imposed by a human-model comparison (since each narrative is long, complex, and requires significant time commitment and detailed evaluation).
-Additionally, from a pool of 20 control participants who underwent the same interview procedure, one control case with no personality disorder and subclinical personality functioning was chosen.
-In total, the evaluation dataset comprised 7 transcriptions of autobiographical interviews: 3 BPD, 3 NPD, and 1 Healthy Control.
+To extract cases for this study, we followed intensity sampling to select information-rich, yet non-extreme cases representing: integrated personality as defined in DSM-5 (HC, $N=1$) and mild, moderate, and severe levels of personality impairment (clinical sample: BPD, $N=3$; NPD, $N=3$).
+The severity variance represented in the sample was meant to capture the breadth of narrative expression within both BPD and NPD.
+The final sample size was set to balance the necessity to represent the heterogeneity of PD symptomatology with the practical constraints imposed by a human evaluation.
+The final database, amounting to 27,641 words, was evaluated for approximately five hours per case on average, thus reaching the maximum volume that avoided overwhelming human evaluators and secured the reliability of their assessments.
 2.2 Model Selection
 We included a total of N=16{N=16} leading models and variants, split into a closed-source ($N=9$) and an open-source group ($N=7$).
-Our selection criteria encompassed: general public interest and availability, presence in recent evaluations on related material fouda2025psychiatrybenchmultitaskbenchmarkllms; hua2025scoping; nguyen-etal-2025-large, access model (licensing), parameter count, country of origin, and reasoning ability.
-The closed-source group included: Gemini 2.5 Pro comanici2025gemini25pushingfrontier, Gemini 3 Pro google2025gemini3, Claude Opus 4.1 anthropic2025claude4, GPT-4o openai2024gpt4ocard, GPT-4.1 openai2025gpt41, and four variants of GPT-5 openai2025gpt5 with all available values of the OpenAI API’s reasoning effort parameter (from minimal to high).
-The open-source group consisted of: Gemma 3 27B gemmateam2025gemma3technicalreport, Llama 3.3 70B grattafiori2024llama3herdmodels, DeepSeek R1 0528 deepseekai2025deepseekr1incentivizingreasoningcapability, two variants of DeepSeek v3.1 Terminus deepseekai2025deepseekv3technicalreport with reasoning either enabled or disabled, and two variants of Qwen 3 32B yang2025qwen3technicalreport, also with and without reasoning.
+Our selection criteria encompassed: public interest and availability, presence in recent evaluations on related material Fouda et al.
+(2025); Hua et al.
+(2025); Nguyen et al.
+(2025), access model (licensing), parameter count, country of origin, and reasoning ability.
+The closed-source group included: Gemini 2.5 Pro Gemini Team (2025), Gemini 3 Pro Google (2025), Claude Opus 4.1 Anthropic (2025), GPT-4o OpenAI (2024), GPT-4.1 OpenAI (2025b), and four variants of GPT-5 OpenAI (2025a) with all available values of the OpenAI API’s reasoning effort parameter (from minimal to high).
+The open-source group consisted of: Gemma 3 27B Gemma Team (2025), Llama 3.3 70B Llama Team (2024), DeepSeek R1 0528 DeepSeek-AI (2025a), two variants of DeepSeek v3.1 Terminus DeepSeek-AI (2025b) with reasoning either enabled or disabled, and two variants of Qwen 3 32B Qwen Team (2025), also with and without reasoning.
 We did not include domain-specific models trained on medical data for two reasons.
-Firstly, the context windows of available medical models evaluated in existing literature fouda2025psychiatrybenchmultitaskbenchmarkllms; hanafi2025comprehensiveevaluationlargelanguage; hua2025scoping; nguyen-etal-2025-large proved to be too small for the full patient testimonies—except for MentalQLM shi2025mentalqlm, which has not been made publicly available.
-Secondly, recent studies indicate that general-purpose models perform similarly to or better than medical models on psychiatric tasks fouda2025psychiatrybenchmultitaskbenchmarkllms.
+Firstly, the context windows of available medical models evaluated in existing literature Fouda et al.
+(2025); Hanafi et al.
+(2025); Hua et al.
+(2025); Nguyen et al.
+(2025) proved to be too small for the full patient testimonies—except for MentalQLM Shi et al.
+(2025), which has not been made publicly available.
+Secondly, recent studies indicate that general-purpose models perform similarly to or better than medical models on psychiatric tasks Fouda et al.
+(2025).
 2.3 Mental Health Experts
 The study sample consisted of $N=6$ highly experienced mental health professionals, comprising three psychiatrists and three psychotherapists, recruited through our professional network.
+They received no compensation for their participation in the study.
 The group was balanced by gender, with three participants identifying as men and three as women; the mean age was 48.5 years (SD = 11.18), with an average of 18.5 years of professional experience.
@@
 They were blinded to the patients’ previous medical records and ground truth diagnoses, and had never previously encountered the patients whose narratives they evaluated.
+Each expert formulated their diagnostic decisions entirely independently.
 2.4 Procedure
@@
 In contrast, both the severity rating and the confidence ratings followed fixed scales: severity was rated on a 0–3 scale (0 = none, 1 = mild, 2 = moderate, 3 = severe), and confidence on a 1–4 scale (1 = guessing, 2 = somewhat confident, 3 = fairly confident, 4 = completely confident).
+Importantly, to reliably evaluate the baseline competence of the models, no model-only in-context learning paradigm was employed in the protocol.
+This is because potential differences in understanding the patients’ data between human participants and models were another point of interest of the evaluation.
+The differences could potentially stem from, e.g., a lack of such priming through example human responses.
 While human participants evaluated each testimony once, models were presented with each case three times to address potential diagnostic inconsistency stemming from their non-deterministic nature.
@@
 2.5.2 Diagnostic Justifications
-Human mental health professionals can arrive at similar diagnostic conclusions through different cognitive and emotional pathways biondi2022clinician.
+Human mental health professionals can arrive at similar diagnostic conclusions through different cognitive and emotional pathways Biondi et al.
+(2022).
 Traces of these mental processes are reflected in the semantic content of the justifications provided during the diagnosis.
@@
 This allowed for a comparison of diagnostic expertise leverage between human and model participants, as well as among different models.
-BAAI/bge-multilingual-gemma2111https://huggingface.co/BAAI/bge-multilingual-gemma2 was chosen as the embedding model owing to its superior performance on Polish-language tasks in the MMTEB benchmark enevoldsen2025mmtebmassivemultilingualtext.
+BAAI/bge-multilingual-gemma2111https://huggingface.co/BAAI/bge-multilingual-gemma2 was chosen as the embedding model owing to its superior performance on Polish-language tasks in the MMTEB benchmark Enevoldsen et al.
+(2025).
 We created a single summary embedding representing the averaged semantic content of justifications for each model separately, as well as a single one representing all human participant data to account for its comparative scarcity.
@@
 The summary embeddings were aggregated into two distinct datasets: a global dataset comprising all agents (both models and humans) to assess the human-AI semantic gap, and a model-only dataset to allow for a finer-grained analysis of inter-model semantic differences.
-To visualize the high-dimensional relationships between these representations, we employed a dual-method dimensionality reduction approach utilizing Multidimensional Scaling (MDS) kruskal1964multidimensional and Uniform Manifold Approximation and Projection (UMAP) mcinnes2020umapuniformmanifoldapproximation.
-This combination was chosen to provide a comprehensive view of the semantic space: MDS was selected for its ability to preserve global metric distances, offering a faithful representation of the relative dissimilarities between models.
-Complementarily, UMAP was utilized for its strength in manifold learning and local structure preservation, which allows for the identification of distinct semantic clusters (neighborhoods).
-MDS was conducted using the scikit-learn [^2] https://github.com/scikit-learn/scikit-learn library JMLR:v12:pedregosa11a.
+Due to the depth-first approach of the study, the summary embeddings constituted an unusually small set of 17 data points, making a faithful visualization of the high-dimensional relationships between these representations harder.
+Instead of the domain standard methods such as UMAP McInnes et al.
+(2020), t-SNE van der Maaten and Hinton (2008), or PCA Pearson (1901), we employed Multidimensional Scaling (MDS) Kruskal (1964) as the dimensionality reduction algorithm.
+Unlike local manifold techniques—unreliable for small sample sizes—or linear projections—which prioritize variance over similarity—MDS was selected for its ability to preserve global metric distances, offering a faithful representation of the relative semantic dissimilarities between the summary embeddings.
+The algorithm was implemented using the scikit-learn [^2] https://github.com/scikit-learn/scikit-learn library Pedregosa et al.
+(2011).
 Before dimensionality reduction, a pairwise cosine distance matrix was computed for the embeddings to ensure that the angular relationships inherent to semantic vector spaces were preserved.
-UMAP was conducted using the umap-learn [^3] https://github.com/lmcinnes/umap library.
 Details on the hyperparameter values used can be found in Appendix F.
@@
 We aggregated all justifications into two distinct corpora—human-written and model-written—and sought to identify lexical features that were statistically overrepresented in one group relative to the other.
-The features were identified with the weighted log-odds ratio with informative Dirichlet prior method originally applied to the problem of detecting lexical polarization in political discourse monroe2008fightin.
+The features were identified with the weighted log-odds ratio with the informative Dirichlet prior method originally applied to the problem of detecting lexical polarization in political discourse Monroe et al.
+(2008).
+The method is well-suited for analyzing small and/or imbalanced text corpora; it estimates the z-score of the log-odds ratio for each n-gram, controlling for variance in word frequency, which specifically addresses the issue of sample size imbalance between the human and model datasets in our use case.
 In the technical implementation, we first applied a text pre-processing pipeline to the two corpora, converting non-Polish non-ASCII characters into ASCII equivalents, as well as removing LLM-characteristic text formatting artifacts, punctuation, digits, and optionally stop words.
-Finally, we applied Monroe et al.’s method through the ConvoKit library chang-etal-2020-convokit.
+Finally, we applied Monroe et al.’s method through the ConvoKit library Chang et al.
+(2020).
 We considered two parameters to adjust in the feature generation process: stop word removal—either enabled or disabled, and the maximum n-gram length—exclusive ranges of 1 to 3 and a combined 1–3 one, ultimately choosing unigrams with the stop words removed as the most informative configuration.
@@
 A comparative analysis of diagnostic performance revealed significant variability among models and mental health professionals, as summarized below.
-Best-performing models outperform the human diagnostic average.
+Top-performing models achieve a higher diagnostic average than human experts.
 The overall human average score was 43.57%, with a categorical score of 3.60 out of 7 (51.43%) and a dimensional score of 2.50 out of 7 (35.71%).
 An exploratory analysis of the participant sub-groups suggested a potential performance gap: psychotherapists scored higher on the categorical task (mean difference: 2.33), while psychiatrists scored slightly higher on the dimensional task (mean difference: 0.33).
+Importantly, inter-rater reliability among human experts was relatively low, reflecting the inherent subjectivity of psychiatric evaluation Di Forti et al.
+(2025) restricted to the text modality Fiquer et al.
+(2013).
+For categorical diagnosis, Fleiss’ Kappa was 0.11, and Krippendorff’s Alpha (nominal) was 0.14, indicating slight agreement; for dimensional severity ratings, Krippendorff’s Alpha (ordinal) was 0.22, indicating fair agreement.
 Model performance, evaluated over 21 trials, was relatively dispersed, with several top-performing models exceeding the human average (see Figure 1).
@@
 Further analysis of model performance indicated that reasoning did not consistently improve diagnostic accuracy.
-Model size appeared to be a factor, with smaller models generally performing worse than larger ones.
-No clear performance gap was observed between closed-source and open-source models.
-BPD overdiagnosis and NPD underdiagnosis drive systematic bias, most pronounced in models.
+Additionally, no clear performance gap was observed between closed-source and open-source models.
+In contrast, model size appeared to be a factor, with smaller models generally performing worse than larger ones.
+However, drawing definitive statements about its influence is difficult, as they would rely exclusively on the open-sourced models in our sample.
+This is further compounded by the fact that model performance being inevitably correlated with model size is (mostly) a general principle of machine learning.
+Diagnostic distributions reveal frequent BPD overdiagnosis and NPD underdiagnosis tendencies, particularly among models.
 An analysis of the $N=378$ total diagnoses (see Appendix C, Table C.1) by both models ($N=336$) and humans ($N=42$) revealed systematic diagnostic bias.
@@
 This suggests a tendency in the GPT model family to favor non-clinical classifications.
-Notably, humans also showed a slight bias toward non-clinical classifications, evidenced by 4 false positives (see Appendix D, Figure D.1), driven by frequent assessments that the testimonies lacked sufficient evidence for a PD diagnosis.
+Notably, humans also demonstrated a slight tendency toward non-clinical classifications, evidenced by 4 false positives (see Appendix D, Figure D.1), driven by frequent assessments that the testimonies lacked sufficient evidence for a PD diagnosis.
 A significant misclassification bias was observed for Avoidant Personality Disorder (AvPD) (see Appendix C, Table C.1).
@@
 AvPD accounted for 16.07% (54 of 336) of model-generated diagnoses, compared to 7.14% (3 of 42) of human diagnoses.
-Given that AvPD was not present in the ground truth answers, this high prevalence points to a systematic tendency to misinterpret symptoms of other conditions as AvPD.
+Given that AvPD was not present in the ground truth answers, this high prevalence points to a tendency to misinterpret symptoms of other conditions as AvPD.
 On the other hand, only one out of 336 diagnoses by models indicated a PD category not present in current diagnostic frameworks, namely Masochistic PD.
@@
 3.2.1 Semantic Embeddings
-Figure 3: MDS and UMAP projections of the semantic embeddings of diagnosis justifications.
-The MDS and UMAP projections of the semantic embeddings of diagnosis justifications are visualized in Figure 3.
-Overall, for this analysis, MDS proved to be a more informative dimensionality-reduction method than UMAP, contributing to most of the observations summarized as follows.
+Figure 3: MDS projections of the semantic embeddings of diagnosis justifications.
+The MDS projections of the semantic embeddings of diagnosis justifications are visualized in Figure 3.
 Models from a single family (but not necessarily the same series) are generally close semantically.
@@
 Llama’s explanations tended to contain incorrect grammar, nonexistent words, or seemingly random ,,artifact” tokens unrelated grammatically and semantically to the rest of the justifications’ linguistic content.
-Interestingly, this did not seem to have significantly impeded Llama’s diagnosis capabilities, as it was the ex-aequo third-best performing model (see Figure 1).
+Interestingly, this did not impede Llama’s diagnosis capabilities, as it was, in fact, the ex-aequo third-best performing model (see Figure 1).
 Models performing weakly in terms of diagnostic performance also exhibit more atypical semantics.
@@
 While human experts performed better on the categorical than the dimensional task, models achieved relatively balanced scores across both, with a performance preference for categorical diagnosis.
-This suggests that the current shift toward a dimensional understanding of personality pathology APA2022; WHO2022 may be overshadowed by the legacy of earlier categorical systems, whose labels remain prevalent in everyday discourse and online corpora.
+This suggests that the current shift toward a dimensional understanding of personality pathology American Psychiatric Association (2022); World Health Organization (2022) may be overshadowed by the legacy of earlier categorical systems, whose labels remain prevalent in everyday discourse and online corpora.
 The prominence of such labels in training corpora may also explain the severe underdiagnosing of NPD by models.
-This could stem from the negative semantic load and social stigma associated with the term ,,narcissist”, combined with the current LLM preference alignment training regimen (i.e., RLHF techniques) rewarding agreeable and non-confrontational behavior, sometimes bordering on (social) sycophancy Cheng2025; sharma2025understandingsycophancylanguagemodels.
-Assigning a stigmatized label to a first-person narrator—essentially ,,calling the user a narcissist”—conflicts with these preferences.
+This could potentially stem from the negative semantic load and social stigma associated with the term ,,narcissist”, combined with the current LLM preference alignment training regimen (i.e., RLHF techniques) rewarding agreeable and non-confrontational behavior, sometimes bordering on (social) sycophancy Cheng et al.
+(2025); Sharma et al.
+(2025).
+We hypothesize that assigning a stigmatized label to a first-person narrator—essentially ,,calling the user a narcissist”—conflicts with these preferences.
 Importantly, this avoidance does not reflect inability, as models showed higher dimensional recall for NPD cases, indicating sensitivity to severity without using value-laden labels.
-A similar aversion to stigmatizing terminology may explain the depathologizing bias observed in the GPT models family, where avoidance of psychiatric labels can downplay genuine symptoms and delay necessary treatment Semigran2015.
+A similar aversion to stigmatizing terminology may explain the depathologizing bias observed in the GPT models family, where avoidance of psychiatric labels can downplay genuine symptoms and delay necessary treatment Semigran et al.
+(2015).
 Conversely, BPD is increasingly treated in the literature as a clinical condition rather than a pejorative label.
-The comparatively extensive body of BPD research relative to other disorders Blashfield2000 likely contributes to the models’ inclination to overdiagnose it.
+The comparatively extensive body of BPD research relative to other disorders Blashfield and Intoccia (2000) likely contributes to the models’ inclination to overdiagnose it.
 Notably, a critical divergence between humans and LLMs was observed in the assessment of confidence.
@@
 As argued by Ulmer et al.
-Ulmer2025, the ability of an AI to express doubt in a ,,human-like” manner is a necessary safety feature to ensure users can trust the uncertainty being conveyed.
+Ulmer et al.
+(2025), the ability of an AI to express doubt in a ,,human-like” manner is a necessary safety feature to ensure users can trust the uncertainty being conveyed.
 Paradoxically, although both humans and models demonstrated higher objective performance on the categorical task, they consistently reported higher subjective certainty regarding their dimensional ratings.
@@
 Llama 3.3 70B was a notable exception, achieving high diagnostic accuracy despite generating justifications riddled with linguistic artifacts, nonexistent words, and poor Polish grammar.
-This can be understood in terms of a multilingual generative-discriminative performance gap, in which Llama 3.3 70B demonstrates sufficient passive understanding of Polish to assign accurate diagnoses, while exhibiting markedly weaker Polish fluency when generating justifications for those decisions.
-Follow-up research could investigate how the performance of this model would change if its diagnostic reasoning were generated before assigning diagnostic scores, rather than as a post-hoc justification.
+This can be understood in terms of a multilingual generative-discriminative performance gap: Llama 3.3 70B demonstrates sufficient understanding of Polish to assign accurate diagnoses, but exhibits markedly weaker Polish generation fluency when creating justifications for those decisions.
+Follow-up research could investigate how the model’s performance would change if its diagnostic reasoning were generated before assigning diagnostic scores, rather than as a post-hoc justification.
 5 Conclusion
-In this study, we provided a first-of-its-kind evaluation of state-of-the-art LLMs on raw, first-person patient data, assessed under a novel expert-informed evaluation protocol and directly compared with human specialists.
-Our findings indicate that current LLMs possess a surprisingly high level of diagnostic competence, largely surpassing mental health professionals in the examined context.
+In this depth-first psychiatric case study, we evaluated state-of-the-art LLMs on raw, first-person patient data, assessed under a novel expert-informed evaluation protocol, and directly compared their performance with human specialists.
+Our observations suggest that current LLMs demonstrate a surprisingly high level of diagnostic competence, achieving higher diagnostic scores than mental health professionals in the examined context.
 At the same time, the models demonstrated susceptibility to bias and overconfidence.
-Taken together, these findings underscore the need for a collaborative human–AI framework, in which the clinical judgment and ethical oversight of experts balance the analytical consistency of models.
+Taken together, these findings underscore the value of a collaborative human–AI framework, in which the clinical judgment and ethical oversight of experts balance the analytical consistency of models.
 Such integration has the potential to mitigate the respective limitations of each agent—model bias and human variability—thereby achieving higher diagnostic validity than either could attain alone.
@@
 This study has several limitations.
-First, humans and LLMs were compared on the textual diagnostic modality only, whereas actual psychiatric assessment is inherently multimodal; clinicians rely on visual cues, affect, prosody, and interactional dynamics that are absent in text transcripts.
-Accordingly, the human experts in our experiment reported difficulties in relying solely on text without seeing the patient.
+First, humans and LLMs were compared on the textual diagnostic modality only, whereas actual psychiatric assessment is inherently multimodal, which could explain the human expert’s comparatively lower performance and low inter-rater reliability in this strictly textual format.
 However, given the rapid post-COVID rise of remote medicine, as well as the more recent surge in AI-driven therapeutic chatbots, the textual modality may gain prominence in the future.
-The second limitation is that the sample comprised only seven narratives.
-Although small, this sample size reflects a deliberate trade-off necessitated by the depth of the data and the substantial cognitive and time demands of human experts.
-The fact that this study was conducted within a specific cultural context may also be considered a limitation.
-This is because the diagnostic approaches of human experts and the narrative styles of patients may not be generalizable to other populations.
-Future research should explore whether these findings remain consistent across different cultural settings.
+Second, although a sample size of seven narratives is notably small by the standards of conventional machine learning evaluations, within the context of this study, it reflects a deliberate tradeoff.
+The limiting factor was the immense cognitive demand of the expert evaluation task.
+Unlike standard clinical vignettes, the cases used in this study consisted of raw, verbatim transcripts of semi-structured autobiographical interviews drawn from a corpus exceeding 200,000 words, often resulting in over 30 pages of chaotic, non-linear narration per patient.
+Properly analyzing just one of these narratives required approximately five hours of expert time, as it involved careful reading, cross-referencing temporal inconsistencies, and synthesizing a coherent clinical picture from a ,,messy” life story.
+Consequently, securing more than 35 hours of active work from practicing, high-level experts proved logistically unfeasible due to time constraints and the significant cognitive toll.
+Expanding the number of narratives would have introduced a severe risk of expert fatigue; the heavy cognitive load required to review such complex qualitative data at scale would have inevitably compromised the depth and quality of the human diagnostic judgments, thereby undermining the depth-first nature of this case study.
+Additionally, since diagnostic approaches of human experts are culture-specific, they may not generalize to other populations.
 Ethical Considerations
@@
 Human participants signed an informed consent form by hand.
-The qualitative interview protocol was approved by the hospital and the Bioethics Committee of Poznan University of Medical Sciences (decision no.
-KB-367/23).
+The qualitative interview protocol was approved by the hospital and the Bioethics Committee of [Anonymized] University of Medical Sciences (decision no.
+[Anonymized]).
 All personal information was anonymized, ensuring full privacy of the patients’ identities.
@@
 Appendix A Clinical Assessment Instruments
-Levels of personality functioning and maladaptive trait domains were assessed using the Level of Personality Functioning—Brief Scale 2.0 (LPFS-BF 2.0) Lakuta2023 and the Personality Inventory for ICD-11 Cieciuch2022, ensuring comparability across individuals.
-LPFS-BF 2.0 scores were interpreted relative to established cut-offs Weekers2023, yielding a dimensional classification from subclinical to extreme impairment (0–48), independent of the categorical BPD/NPD labels.
-Appendix B Model Prompts
+Levels of personality functioning and maladaptive trait domains were assessed using the Level of Personality Functioning—Brief Scale 2.0 (LPFS-BF 2.0) Łakuta et al.
+(2023) and the Personality Inventory for ICD-11 Cieciuch et al.
+(2022), ensuring comparability across individuals.
+LPFS-BF 2.0 scores were interpreted relative to established cut-offs Weekers et al.
+(2023), yielding a dimensional classification from subclinical to extreme impairment (0–48), independent of the categorical BPD/NPD labels.
+Appendix B Prompts
 Original Polish prompt used in the study Jesteś doświadczonym specjalistą zdrowia psychicznego.
@@
 random_state 42
-UMAP n_components 2
-n_neighbors 4
-min_dist 0.1
-random_state 42
-Table F.2: Dimensionality reduction algorithms hyperparameter values.
-Hyperparameters not present in the table retained the default values set in the algorithm implementations.
-The n_neighbors parameter for UMAP, which controls the balance between local and global structure, was set to a low value to accommodate the small dataset size of $N=17/16$ (with and without the human embedding).
-This setting prioritizes the detection of latent local substructures and family-specific clusters within the models.
+Table F.2: Dimensionality reduction algorithm hyperparameter values.
+Hyperparameters not present in the table retained the default values set in the algorithm implementation.
 Appendix G Semantic Justification Embedding Creation Process